# A Random Matrix Theory of Pauli Tomography

### v1

@@ III The GUE in the Pauli Basis
In this section we are concerned with relationships between probability measures on subspaces of the space $\mathrm{Mat}_{D}(\mathbb{C})$ of $D\times D$ matrices over $\mathbb{C}$, and probability measures on $\mathbb{R}^{D^{2}}$. The central object of interest is the Gaussian Unitary Ensemble (GUE), a probability measure on the $D^{2}$ dimensional real subspace of Hermitian matrices $\mathcal{H}_{D}\subset\mathrm{Mat}_{D}(\mathbb{C})$ with density given by independent Gaussians over its independent entries (i.e. those elements not fixed by enforcing hermiticity), each with variance $\sigma^{2}$ [43]:
$\mathrm{d}\mu(H)$ $\propto\exp\left(-\frac{D}{2\sigma^{2}}\left[\sum_{j\leq k}|H_{jk}^{(\mathrm{r})}|^{2}+\sum_{j<k}|H_{jk}^{(\mathrm{i})}|^{2}\right]\right)\mathrm{d}H$ (4)
$=\exp\left(-\frac{D}{2\sigma^{2}}\mathrm{Tr}(H^{2})\right)\mathrm{d}H,$ (5)
@@
with radius $R=2\sigma$.
@@
where $\mathrm{d}a$ is the Lebesgue measure on $\mathbb{R}^{D^{2}}$. Here, $d\mu_{\Sigma}$ corresponds to sampling from the multivariate Gaussian distribution $\mathcal{N}(0,\Sigma)$ with means 00 and covariance matrix $\Sigma$. Such a random vector $\vec{a}$ naturally arises in tomographic settings (see Section II); and the map Eq. 8 can be understood as a tomographic reconstruction protocol given some experimentally realizable IC-basis.
@@
which is identical in form to the GUE measure of Eq. 10 with rescaled variances $\sigma^{2}=vDC$. ∎
@@
Proof. The proof proceeds identically to that of Theorem 2; with the notable substitution of the SIC-POVM effect definition $E_{j}=\Pi_{j}/D$ where $P_{j}$ is a projector, and the SIC-POVM overlap definition $\mathrm{Tr}(\Pi_{j}\Pi_{k})=D(\delta_{jk}+1)/(D+1)$. Proceeding identically to the calculation of Eq. 17, we find the SIC-POVM quadratic form
@@ IV Reconstruction error as a likely sampling of the GUE
$\mathbb{P}\left[\left|\int f\mathrm{d}\mu_{y}-\int f\mathrm{d}\mu_{g}\right|^{2}\geq a||f||_{\mathrm{Lip}}^{2}\right]\leq\frac{D}{a\overline{v}}\mathrm{var}[v],$ (21)
@@
We now turn our attention to the evaluation of $\mathbb{E}[\|\mathcal{A}_{E}(\vec{y})-\mathcal{A}_{E}(\vec{g})\|_{F}^{2}]$. We proceed by fixing a form for $\vec{g}$ based on $\vec{y}$ and computing the expectation value directly; first rewrite the elements of $\vec{y}$ as $y_{j}=z_{j}\sqrt{v_{j}}$ in terms of the dummy statistic $z_{j}\sim\mathcal{N}(0,1)\sqrt{v_{j}}$. We can then introduce a corresponding vector $\vec{g}$ of values $g_{j}=z_{j}\sqrt{\overline{v}}$, such that $(y_{j}-g_{j})/z_{j}=\eta_{j}=\sqrt{v_{j}}-\sqrt{\overline{v}}$. By Theorem 2, $\mathcal{A}_{E}(\vec{g})$ is a GUE sampled matrix. Using Eq. 8 we can explicitly construct the Frobenius norm as follows:
@@ IV.2 QWC Tomographic Protocol
To evaluate the first term on the r.h.s., note that there are (${10\choose k}3^{k}$ Pauli strings with weight $k$, and hence:
@@ V.1 Analytic Results for Quantum State Tomography Errors
$\|\hat{\rho}-\hat{\rho}^{\prime}\|_{\mathrm{Tr}}=\|\Delta\hat{\rho}\|_{\mathrm{Tr}}=\sum_{j}|\lambda_{j}(\Delta\hat{\rho})|,$ (57)
@@
[…]
@@
[…]
@@
±=±$\mathcal{I}_{\pm}=\pm\int^{R}_{0}\mathrm{d}\lambda_{1}\mathcal{P}\left[\int^{R}_{0}\mathrm{d}\lambda_{2}\frac{\lambda_{1}}{\lambda_{1}\mp\lambda_{2}}\sqrt{\frac{R^{2}-\lambda_{2}^{2}}{R^{2}-\lambda_{1}^{2}}}\right]$ (65)
respectively. We can explicitly compute ±$\mathcal{I}_{\pm}=R^{2}(4\pm\pi^{2})/8$ by factorizing the integrand into a form to which generic integral formulae apply, and taking limits appropriately to resolve the principal value. Direct substitution into Eq. 64 yields the variance
@@ V.2 Independent and Correlated Coefficients
In Fig. 2(a) we show the eigenvalue distributions of a single realization of $\Delta\hat{\rho}$ for both states respectively, and have superimposed the Wigner semicircle distribution of radius $R=2/\sqrt{S}$ (using the naïve protocol’s value for $\overline{v}$) on Fig. 2(a) ; and can clearly see agreement between all three sets of data. This suggests an even stronger statement than those made in Section IV; wherein for the case of naïve QST in the Pauli IC-basis, the ensemble of $\Delta\hat{\rho}$ itself converges to the GUE. Panels (b) and (c) show the numerically determined rescaled expectation values $\mathbb{E}[\|\Delta\hat{\rho}\|_{\mathrm{Tr}}]/2^{N}$ for the GHZ and identity states respectively, and panels (d) and (e) show the corresponding variances $\mathrm{var}[\|\Delta\hat{\rho}\|_{\mathrm{Tr}}]$. We have taken 100 sample realizations of $\Delta\hat{\rho}$ for every data point, and have superimposed the (rescaled) analytic values of Eq. 67 and Eq. 68. We find strong agreement at all system sizes; with slight deviation at $N=2$ for the GHZ state due to the fact that our main results (and thus derived expected values and variances) hold at leading order in $N$. The reason that the identity state does not exhibit this system-size dependence, and the reason that it shows almost exact agreement with our analytic prediction, is because all the Pauli coefficients (excluding the trivial all-identity Pauli string) are zero, and thus induce an almost-uniform diagonal $\Sigma$ allowing us to apply Eq. 18 directly.
@@ V.4 Comparison to the Maximum Likelihood Estimator
[…]
@@
The first term in the absolute value sign on the left is of order $\left(10/3\right)^{N}/\sqrt{S}$, while the second is bounded by a term of order $\left(5/6\right)^{N/2}/\sqrt{S}$. From this lower bound, and the fact that $\hat{\Pi}$ is contractive, we must have:
$\mathbb{E}[\|\hat{\Pi}\hat{\rho}^{\prime}-\hat{\rho}\|_{\mathrm{Tr}}]\sim\Theta\left(\frac{1}{\sqrt{S}}\left(\frac{10}{3}\right)^{N}\right),$ (86)
@@
We believe but are unable to show in this work that any rephysicalization procedure necessarily cannot improve the sample procedure of tomographic reconstruction, since it will involve mapping an object from $\mathrm{hut}_{D}$ to a convex subset of linear codimension 00. We offer the above result in support of this claim, with the view that Eq. 86 holds for any single-copy, non-adaptive QST estimator for the a priori $\hat{\rho}$.
@@ V.5 Other Directions
We have demonstrated in Theorem 4, Theorem 5, Theorem 6 that the discussed tomographic protocols yield GUE results at leading order for Lipschitz functions over the one and two point empirical spectral measures (and as an easy generalisation, any $k$-point empirical spectral measure where $k$ is fixed for growing $N$); whether this correspondence holds for other functions of the spectrum of $\Delta\hat{\rho}$ and what functions would be useful - may be fruitful questions.

### v3

@@ III The GUE in the Pauli Basis
In this section we are concerned with relationships between probability measures on subspaces of the space $\mathrm{Mat}_{D}(\mathbb{C})$ of $D\times D$ matrices over $\mathbb{C}$, and probability measures on $\mathbb{R}^{D^{2}}$. The central object of interest is the Gaussian Unitary Ensemble (GUE), a probability measure on the $D^{2}$ dimensional real subspace of Hermitian matrices $\mathcal{H}_{D}\subset\mathrm{Mat}_{D}(\mathbb{C})$ with density given by independent Gaussians over its independent entries (i.e. those elements not fixed by enforcing hermiticity), each with total variance $\sigma^{2}$ [43]:
$\mathrm{d}\mu(H)$ $\propto\exp\left(-\sum_{j}\frac{H_{jj}^{2}}{2\sigma^{2}}-\sum_{j<k}\frac{[H_{jk}^{(r)}]^{2}}{\sigma^{2}}-\sum_{j<k}\frac{[H_{jk}^{(i)}]^{2}}{\sigma^{2}}\right)\mathrm{d}H$ (4)
$=\exp\left(-\frac{1}{2\sigma^{2}}\mathrm{Tr}(H^{2})\right)\mathrm{d}H,$ (5)
@@
with radius $R=2\sigma\sqrt{D}$.
@@
where $\mathrm{d}a$ is the Lebesgue measure on $\mathbb{R}^{D^{2}}$. Here, $d\mu_{\Sigma}$ corresponds to sampling from the multivariate Gaussian distribution $\mathcal{N}(0,\Sigma)$ with means 0 and covariance matrix $\Sigma$. Such a random vector $\vec{a}$ naturally arises in tomographic settings (see Section II); and the map Eq. 8 can be understood as a tomographic reconstruction protocol given some experimentally realizable IC-basis.
@@
which is identical in form to the GUE measure of Eq. 10 with rescaled variances $\sigma^{2}=vC$. ∎
@@
Proof. The proof proceeds identically to that of Theorem 2; with the notable substitution of the SIC-POVM effect definition $E_{j}=\Pi_{j}/D$ where $\Pi_{j}$ is a projector, and the SIC-POVM overlap definition $\mathrm{Tr}(\Pi_{j}\Pi_{k})=D(\delta_{jk}+1)/(D+1)$. Proceeding identically to the calculation of Eq. 17, we find the SIC-POVM quadratic form
@@ IV Reconstruction error as a likely sampling of the GUE
[…]
@@
We now turn our attention to the evaluation of $\mathbb{E}[\|\mathcal{A}_{E}(\vec{y})-\mathcal{A}_{E}(\vec{g})\|_{F}^{2}]$. We proceed by fixing a form for $\vec{g}$ based on $\vec{y}$ and computing the expectation value directly; first rewrite the elements of $\vec{y}$ as $y_{j}=z_{j}\sqrt{v_{j}}$ in terms of the dummy statistic $z_{j}\sim\mathcal{N}(0,1)$. We can then introduce a corresponding vector $\vec{g}$ of values $g_{j}=z_{j}\sqrt{\overline{v}}$, such that $(y_{j}-g_{j})/z_{j}=\eta_{j}=\sqrt{v_{j}}-\sqrt{\overline{v}}$. By Theorem 2, $\mathcal{A}_{E}(\vec{g})$ is a GUE sampled matrix. Using Eq. 8 we can explicitly construct the Frobenius norm as follows:
@@ IV.2 QWC Tomographic Protocol
To evaluate the first term on the r.h.s., note that there are ${10\choose k}3^{k}$ Pauli strings with weight $k$, and hence:
@@ V.1 Analytic Results for Quantum State Tomography Errors
[…]
@@
$\mathbb{E}[\|\Delta\hat{\rho}\|_{\mathrm{Tr}}]=D\int_{-R_{S}}^{R_{S}}|\lambda|W_{R}(\lambda)d\lambda.$ (61)
@@
$\mathbb{E}[\|\Delta\hat{\rho}\|_{\mathrm{Tr}}]=2D\int_{0}^{R_{S}}\lambda W_{R}(\lambda)d\lambda=\frac{4D}{3\pi}R,$ (62)
@@
$\mathcal{I}_{\pm}=\pm\int^{R}_{0}\mathrm{d}\lambda_{1}\mathcal{P}\left[\int^{R}_{0}\mathrm{d}\lambda_{2}\frac{\lambda_{1}}{\lambda_{1}\mp\lambda_{2}}\sqrt{\frac{R^{2}-\lambda_{2}^{2}}{R^{2}-\lambda_{1}^{2}}}\right]$ (65)
respectively. We can explicitly compute $\mathcal{I}_{\pm}=R^{2}(4\pm\pi^{2})/8$ by factorizing the integrand into a form to which generic integral formulae apply, and taking limits appropriately to resolve the principal value. Direct substitution into Eq. 64 yields the variance
@@ V.2 Independent and Correlated Coefficients
In Fig. 2(a) we show the eigenvalue distributions of a single realization of $\Delta\hat{\rho}$ for both states respectively, and have superimposed the Wigner semicircle distribution of radius $R=2/\sqrt{S}$ (using the naïve protocol’s value for $\overline{v}$) on Fig. 2(a) ; and can clearly see agreement between all three sets of data. This suggests an even stronger statement than those made in Section IV; wherein for the case of naïve QST in the Pauli IC-basis, the ensemble of $\Delta\hat{\rho}$ itself converges to the GUE. Panels (b) and (c) show the numerically determined rescaled expectation values $\mathbb{E}[\|\Delta\hat{\rho}\|_{\mathrm{Tr}}]/2^{N}$ for the GHZ and identity states respectively, and panels (d) and (e) show the corresponding variances $\mathrm{var}[\|\Delta\hat{\rho}\|_{\mathrm{Tr}}]$. We have taken 100 sample realizations of $\Delta\hat{\rho}$ for every data point, and have superimposed the (rescaled) analytic values of Eq. 67 and Eq. 68. We find strong agreement at all system sizes; with slight deviation at $N=2$ for the GHZ state due to the fact that our main results (and thus derived expected values and variances) hold at leading order in $N$. The reason that the identity state does not exhibit this system-size dependence, and the reason that it shows almost exact agreement with our analytic prediction, is because all the Pauli coefficients (excluding the trivial all-identity Pauli string) are zero, and thus induce an almost-uniform diagonal $\Sigma$ - allowing us to apply Eq. 18 directly.
@@ V.4 Comparison to the Maximum Likelihood Estimator
$\|\hat{\Pi}\hat{\rho}^{\prime}-\hat{\rho}^{\prime}\|_{\mathrm{Tr}}=\left|\int^{R}_{0}\frac{d\mu}{d\lambda}f(\lambda)\,d\lambda\right|,$ (82)
@@
The first term in the absolute value sign on the left is of order $\left(10/3\right)^{N/2}/\sqrt{S}$, while the second is bounded by a term of order $\left(5/6\right)^{N/2}/\sqrt{S}$. From this lower bound, and the fact that $\hat{\Pi}$ is contractive, we must have:
$\mathbb{E}[\|\hat{\Pi}\hat{\rho}^{\prime}-\hat{\rho}\|_{\mathrm{Tr}}]\sim\Theta\left(\frac{1}{\sqrt{S}}\left(\frac{10}{3}\right)^{N/2}\right),$ (86)
@@
We believe but are unable to show in this work that any rephysicalization procedure necessarily cannot improve the sample procedure of tomographic reconstruction, since it will involve mapping an object from $\mathrm{hut}_{D}$ to a convex subset of linear codimension 0. We offer the above result in support of this claim, with the view that Eq. 86 holds for any single-copy, non-adaptive QST estimator for the a priori $\hat{\rho}$.
@@ V.5 Other Directions
We have demonstrated in Theorem 4, Theorem 5, Theorem 6 that the discussed tomographic protocols yield GUE results at leading order for Lipschitz functions over the one and two point empirical spectral measures (and as an easy generalisation, any $k$-point empirical spectral measure where $k$ is fixed for growing $N$); whether this correspondence holds for other functions of the spectrum of $\Delta\hat{\rho} -$ and what functions would be useful - may be fruitful questions.
@@ VII Acknowledgements
VIII Author Contributions
A.N.-K. and N.K. both conceptualised this work, carried out numerical experiments, and wrote the manuscript. N.K. developed the bulk of the mathematical formalisation. All authors contributed to reviewing and editing the manuscript.